# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v2

@@ V-C1 State-based Observations
Next, we evaluate the performance of the policies for different NFE in the testing phase. For RFMP and SRFMP, we use the $50$-epoch models for Lift, Can, and Square, and the $100$-epoch model for Tool Hang. DP is further trained for a total of $300$ epochs and we select the model at the best validation epoch. The results are reported in Table VIII. RFMP and SRFMP outperform DP for all task and all NFE, even though DP was trained for more epochs. Moreover, we observe that RFMP and SRFMP are generally more robust to low NFE than DP. They achieve $100$% success rate at almost all NFE values for the easier Lift and Can tasks, while DP’s performance drastically drops for $1$ and $2$ NFE. We observe a similar trend for the Square task, where the performance of RFMP and SRFMP slightly improves when increasing the NFE. The performance of all models drops for Tool Hang, which is the most complex of the four considered tasks. In this case, the performance of RFMP and SRFMP is limited for low NFE values and improves for higher NFE. DP performs poorly for all considered NFE values. Table IX reports the jerkiness as a measure of the smoothness of the trajectories generated by the different policies. We observe that RFMP and SRFMP produce arguably smoother trajectories than DP for low NFE, as indicated by the lower jerkiness values. The smoothness of the trajectories becomes comparable for higher NFE. In summary, both RFMP and SRFMP achieve high success rates and smooth action predictions with low NFE, enabling faster inference without compromising task completion.
@@ V-D Franka Kitchen Benchmark
Each model is trained for $500$ epochs with a batch size of $32$, and evaluated using the best validation checkpoint over $50$ test episodes with randomized initial states. The performance of different policies is reported in Table XII. SRFMP exhibits consistently high performance across all NFE settings, indicating strong robustness, and significantly outperforms RFMP and DP. The low success rate of RFMP and DP is due to the fact that their trajectories often exhibit local inaccuracies, leading to failing at least one of the four subtasks. This is then reported as a failure for the long-horizon task, leading to low success rates overall. In contrast, the trajectories produced by SRFMP are more precise, leading to the successful completion of all subtasks.
@@ V-E1 Experimental Setup
Figure 9 shows our experimental setup. The tasks are performed on a Franka Emika Panda robot arm. We collect the demonstrations via a teleoperation system consisting of two robot twins. Demonstrations are collected by an expert who guides the source robot. The target robot then reads the end-effector pose of the source robot and reproduces it via a Cartesian impedance controller. The demonstration data were smoothed to mitigate the noise arising from the teleoperation setup. Each observation ${\bm{o}}^{s}$ is composed of the end-effector position, and of the image embedding obtained from the ResNet vision backbone that processes the images from an over-the-shoulder camera. The policies are trained to generate $8$-dimensional actions composed of the position, orientation, and gripper state lying on the product of manifold $\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}$.
Figure 9: Robotic experimental setup consisting of $2$ Franka Emika Panda robot arms for teleoperation and an over-the-shoulder camera (Realsense d435). The left arm is follower robot, while the right one acts as the leader. During the teaching phase, a human expert controls the leader arm to send the desired reference to the follower robot via ROS. During testing, only the follower arm is operational.
@@ V-E2 Pick & Place
The goal of this task is to test the ability of RFMP and SRFMP to learn Euclidean policies in real-world settings. The task consists of approaching and picking up a white mug, and to then place it on a pink plate, as shown in Figure 10. Note that the end effector of the robot points downwards during the entire task, so that its orientation remains almost constant.
@@ V-F Experiments Summary
We conducted comprehensive evaluations across diverse benchmarks in simulation, namely the Euclidean and Sphere Push-T tasks, the Robomimic benchmark under both state- and vision-based observation, the long-horizon Franka Kitchen benchmark, as well as two real-world manipulation tasks. These tasks display various action spaces with different types of geometric constraints. While the Euclidean Push-T and Robomimic tasks are defined on standard Euclidean spaces — a special type of Riemannian manifolds — the Sphere Push-T task involves an action space constrained to the hypersphere $\mathbb{S}^{2}$, and both the Franka Kitchen and real-world experiments operate on product Euclidean and sphere manifolds, accounting for position, orientation, and gripper state. Through their Riemannian formulation, RFMP and SRFMP are designed to naturally handle Euclidean and non-Euclidean action spaces, thus satisfying constraints such as the unit-norm of quaternions. It is important to emphasize that the Riemannian formulation is necessary not only to handle geometric constraints but also to guarantee stability in SRFMP. Naive approximations, e.g., unit-norm normalization as a post-processing step for quaternions, would break the stability guarantees in SRFMP.

### v3

@@ V-C1 State-based Observations
Next, we evaluate the performance of the policies for different NFE in the testing phase. For RFMP and SRFMP, we use the $50$-epoch models for Lift, Can, and Square, and the $100$-epoch models for Tool Hang. DP is further trained for a total of $300$ epochs and we select the model at the best validation epoch. The results are reported in Table VIII. RFMP and SRFMP outperform DP for all tasks and all NFE, even though DP was trained for more epochs. Moreover, we observe that RFMP and SRFMP are generally more robust to low NFE than DP. They achieve 100100% success rate at almost all NFE values for the easier Lift and Can tasks, while DP’s performance drastically drops for $1$ and $2$ NFE. We observe a similar trend for the Square task, where the performance of RFMP and SRFMP slightly improves when increasing the NFE. The performance of all models drops for Tool Hang, which is the most complex of the four considered tasks. In this case, the performance of RFMP and SRFMP is limited for low NFE values and improves for higher NFE. DP performs poorly for all considered NFE values. Table IX reports the jerkiness as a measure of the smoothness of the trajectories generated by the different policies. We observe that RFMP and SRFMP produce arguably smoother trajectories than DP for low NFE, as indicated by the lower jerkiness values. The smoothness of the trajectories becomes comparable for higher NFE. In summary, both RFMP and SRFMP achieve high success rates and smooth action predictions with low NFE, enabling faster inference without compromising task completion.
@@ V-D Franka Kitchen Benchmark
Each model is trained for $500$ epochs with a batch size of $32$, and evaluated using the best validation checkpoint over $50$ test episodes with randomized initial states. The performance of different policies is reported in Table XII. SRFMP exhibits consistently high performance across all NFE, indicating strong robustness, and significantly outperforms RFMP and DP. The low success rate of RFMP and DP is due to the fact that their trajectories often exhibit local inaccuracies, leading to failing at least one subtask. This is then reported as a failure for the long-horizon task, leading to low success rates overall. In contrast, the trajectories produced by SRFMP are more precise, leading to the successful completion of all subtasks.
@@ V-E1 Experimental Setup
Figures 9-11 show our experimental setup. The tasks are performed on a Franka Emika Panda robot arm. We collect the demonstrations via a teleoperation system made of two robot twins. Demonstrations are collected by an expert guiding the source robot. The target robot reads the end-effector pose of the source robot and reproduces it via a Cartesian impedance controller. The demonstration data was smoothed to mitigate the noise arising from the teleoperation setup. Each observation ${\bm{o}}^{s}$ is composed of the end-effector position and of the image embedding obtained from the ResNet vision backbone that processes the images from an over-the-shoulder camera. The policies are trained to generate $8$-dimensional actions composed of the position, orientation, and gripper state lying on the product of manifolds $\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}$.
Figure 9: Robotic experimental setup consisting of two Franka Emika Panda robot arms and an over-the-shoulder camera (Realsense d435). The left arm is the target robot, while the right one acts as the source. During the teaching phase, a human expert controls the source arm to teleoperate the target robot. During testing, only the target arm is operational.
@@ V-E2 Pick & Place
The goal of this task is to test the ability of RFMP and SRFMP to learn Euclidean policies in real-world settings. The task consists of approaching and picking up a white mug, and place it on a pink plate, as shown in Figure 10. Note that the robot end-effector points downwards during the entire task, so that its orientation remains almost constant.
@@ V-F Experiments Summary
We conducted comprehensive evaluations across diverse benchmarks in simulation, namely the Euclidean and Sphere Push-T tasks, the Robomimic benchmark under both state- and vision-based observation, the long-horizon Franka Kitchen benchmark, as well as two real-world manipulation tasks. These tasks display various action spaces with different types of geometric constraints. While the Euclidean Push-T and Robomimic tasks are defined on standard Euclidean spaces — a special type of Riemannian manifolds — the Sphere Push-T task involves an action space constrained to the hypersphere $\mathcal{S}^{2}$, and both the Franka Kitchen and real-world experiments operate on product Euclidean and sphere manifolds, accounting for position, orientation, and gripper state. Through their Riemannian formulation, RFMP and SRFMP are designed to naturally handle Euclidean and non-Euclidean action spaces, thus satisfying constraints such as the unit-norm of quaternions. It is important to emphasize that the Riemannian formulation is necessary not only to handle geometric constraints but also to guarantee stability in SRFMP. Naive approximations, e.g., unit-norm normalization as a post-processing step for quaternions, would break the stability guarantees in SRFMP.